OpenStudy (anonymous):

What is the maximum height, in metres reached by the projectile: h = -1/2 gt^2 + vot + ho?

OpenStudy (anonymous):

g = 9.8 m/s^2 t seconds vo = m/s initial velocity ho = m, initial height of the object. Vi = initial velocity of 34.4 m/s and is launched 2.1 m above the ground.